OpenStudy (anonymous):

Find x and RS if S is between R and T. RS= 6x , ST= 12 < and RT= 72

hero (hero):

RS + ST = RT 6x + 12 = 72 6x = 72 - 12 6x = 60 x = 60/6 x = 10

hero (hero):